# LtU-ILI: An All-in-One Framework for Implicit Inference in Astrophysics and Cosmology

### v1

@@ 2.2 Neural Density Estimation
In modern applications, implicit inference is made tractable by using machine learning models to emulate conditional probability distributions, a procedure called Neural Density Estimation (Papamakarios(2019)). As an example, consider a situation wherein we seek to train a mixture density network (MDN; Bishop(1994)) to directly model the conditional distribution $\mathcal{P}(\bm{v}|\bm{u})$ using our training set of data-parameter pairs $(\bm{x},\bm{\theta})\sim\mathcal{D}_{\rm train}$. For this subsection only, we have $(\bm{u},\bm{v})=(\bm{x},\bm{\theta})$ or $(\bm{u},\bm{v})=(\bm{x},\bm{\theta})$ interchangeably. The goal is to build a neural architecture $q_{\bm{w}}(\bm{v}|\bm{u})$ with weights $\bm{w}$ that outputs a probability distribution over $\bm{v}$ which targets the conditional probability $\mathcal{P}(\bm{v}|\bm{u})$. An MDN outputs probability distributions by allowing the neural network to specify the statistical parameters of a mixture density distribution. In the independent Gaussian case, this means the neural network outputs a mean and a variance for each component of the mixture density distribution, i.e.,
@@
where the expectation $\mathbb{E}_{\mathcal{D}_{\rm train}}$ is taken over all data-parameter pairs in the training set $\sim\mathcal{D}_{\rm train}$. For the case of the the posterior distribution ($\bm{v}=\bm{\theta},\bm{u}=\bm{x}$), if the training data is sufficiently diverse and if the MDN is sufficiently flexible, then the minimization of $\mathcal{L}_{\rm MDN}$ over the network weights $\bm{w}$ will converge $q_{\bm{w}}(\bm{\theta}|\bm{x})$ to the true posterior $\mathcal{P}(\bm{\theta}|\bm{x})$ (Papamakarios and Murray(2016)).
The MDN is just one example of an ever-increasing variety of neural density estimators in the literature. Another very popular class of methods is normalizing flows (Papamakarios et al.(2021)Papamakarios, Nalisnick, Rezende, Mohamed, and Lakshminarayanan), which output a flexible conditional distribution defined via learnable, invertible transformations of a base Gaussian distribution. Normalizing flows are often better equipped for handling cases wherein the target distribution is non-Gaussian, but low-dimensional and uni-modal. For any type of neural density estimator, the training procedure always involves minimizing the negative log-probability of training data, as in Equation 3.
@@ 2.2.1 Neural Posterior Estimation (NPE)
The most straightforward method to do this is to train neural networks to directly emulate the posterior distribution, also known as Neural Posterior Estimation (NPE; Papamakarios and Murray(2016); Greenberg et al.(2019)Greenberg, Nonnenmacher, and Macke). As in the MDN example in the previous section, the loss function for NPE is simply the negative log-posterior of our learned neural density estimator $\hat{\mathcal{P}}(\bm{\theta}|\bm{x})$, and can be trained using a loss function like that of Papamakarios and Murray(2016):
@@
where our neural posterior $\hat{\mathcal{P}}(\bm{\theta}_{i}|\bm{x}_{i})$ is decomposed into a neural network output $q_{\bm{w}}(\bm{x},\bm{\theta})$ and a weighting factor taken as the ratio of the assumed prior $p(\bm{\theta})$ to the proposal prior $\tilde{p}(\bm{\theta})$. The proposal prior is representative of the distribution of $\bm{\theta}$ in the training dataset $\mathcal{D}_{\rm train}$ and is often taken to be equal to the assumed prior. Note, that the normalization factor $\log p(\bm{\theta})/\tilde{p}(\bm{\theta})$ cancels out when optimizing $\mathcal{L}_{\rm NPE}$ over $\bm{w}$, but is still required for constructing the posterior estimator $\hat{\mathcal{P}}(\bm{\theta}|\bm{x})$. This method has the advantage that it allows for quick evaluation and sampling of the full posterior at inference time. However, the disadvantage is that it requires knowledge of the analytic form of the proposal prior $\tilde{p}(\bm{\theta})$ from which training data was sampled, which may not be accessible for astrophysical training data (e.g., cluster properties in cosmological simulations). Recently, Vasist et al.(2023)Vasist, Rozet, Absil, Mollière, Nasedkin, and Louppe used an NPE strategy to constrain exoplanetary atmospheric models. NPE has also become a known tool in Gravitational-Wave astronomy where analytical Compact Binary Coalescent (CBC) waveforms play the role of the forward model (e.g., with DINGO from Dax et al.(2021a)Dax, Green, Gair, Macke, Buonanno, and Schölkopf, and the constraints on GW150914 by Crisostomi et al.(2023)Crisostomi, Dey, Barausse, and Trotta, among many others).
@@ 2.2.2 Neural Likelihood Estimation (NLE)
An alternative method which circumvents the issue of fixed priors is to only fit for the likelihood, $\mathcal{P}\left(\bm{x}|\bm{\theta}\right)$, via Neural Likelihood Estimation (NLE; Alsing et al.(2018)Alsing, Wandelt, and Feeney; Alsing et al.(2019)Alsing, Charnock, Feeney, and Wandelt; Papamakarios et al.(2019)Papamakarios, Sterratt, and Murray). The loss function for NLE notably does not include the assumed or proposal prior, and simply maximizes the global log-likelihood,
@@
wherein we emphasize the swapped arguments of $q$ as compared to Equation 4. Multiplying a learned likelihood with an assumed prior results in a proxy which is proportional to the posterior, i.e., $\hat{\mathcal{P}}(\bm{\theta}|\bm{x})\propto q_{\bm{w}}(\bm{x}|\bm{\theta})p(\bm{\theta})$. Generating samples from the posterior is then simply a matter of using this proxy in running Markov chain Monte Carlo (MCMC; Robert et al.(1999)Robert, Casella, and Casella) or Variational Inference (Blei et al.(2017)Blei, Kucukelbir, and McAuliffe). For example, in Metropolis-Hastings MCMC (Robert et al.(2004)Robert, Casella, Robert, and Casella), the learned likelihood and assumed prior can be used to calculate the acceptance probability of a transition $\bm{\theta}_{t}\rightarrow\bm{\theta}^{\prime}$ without knowledge of the full normalized posterior, i.e.,
@@
where $q(\bm{\theta}^{\prime}|\bm{\theta}_{t})$ is the proposal mechanism. Note, the process of obtaining a posterior is identical to the explicit likelihood-based approach, but instead of an analytic likelihood function, NLE has essentially a “black box” function instead. The advantage of this method is that one only needs to perform training once for a given model, at the cost of additional sampling once data becomes available. For instance, Jeffrey et al.(2021)Jeffrey, Alsing, and Lanusse fit a density model for the likelihood of compressed summaries from DES SV weak lensing maps. The Madminer tool (Brehmer et al.(2020)Brehmer, Kling, Espejo, and Cranmer) for particle physics notably allows for NLE using ILI tools.
@@ 2.2.3 Neural Ratio Estimation (NRE)
As in NLE, Neural Ratio Estimation (NRE; Hermans et al.(2020)Hermans, Begy, and Louppe) targets a proxy that is proportional to the posterior. Instead of outputting a conditional density estimate for the likelihood, NRE models instead target the likelihood ratio,
@@
equal to the ratio of the likelihood at a given point in parameter space to the likelihood at a reference point, $\bm{\theta}_{0}$. This formulation is convenient because it can be folded into an acceptance ratio like Equation 6 and also be cast as a classification problem (Cranmer et al.(2015)Cranmer, Pavez, and Louppe). The output of a classification network, $d_{\bm{w}}(\bm{x},\bm{\theta})\in[0,1]$, can be considered equivalent to a likelihood ratio as
@@
wherein we associate a positive classification with assigning a data $\bm{x}$ to the correct $\bm{\theta}$, and a negative classification with those assigning data to a $\bm{\theta}^{\prime}$ sampled from the prior. The main benefit of NRE over NPE and NLE is that one need not specify an approximating distribution such as MDNs or normalizing flows to emulate a posterior. The complexity is purely limited in terms of the depth and design of the neural architecture. Among others, Cole et al.(2022)Cole, Miller, Witte, Cai, Grootes, Nattino, and Weniger and Karchev et al.(2023)Karchev, Trotta, and Weniger applied a variation of NRE (namely Truncated Marginal NRE, Miller et al.(2021)Miller, Cole, Forré, Louppe, and Weniger) to CMB data and Supernova data, respectively. Additionally, Delaunoy et al.(2020)Delaunoy, Wehenkel, Hinderer, Nissanke, Weniger, Williamson, and Louppe and Bhardwaj et al.(2023)Bhardwaj, Alvey, Miller, Nissanke, and Weniger use NRE for Gravitational Wave parameter inference.
2.2.4 Choosing an Estimator
@@
The optimal choice of NPE, NLE, or NRE is highly dependent on the properties on the underlying problem, the dimensionalities of $\bm{x}$ and $\bm{\theta}$, and the intended application. First, the ease of learning the shape of a likelihood or posterior distribution can vary dramatically between different physical problems. For example, the shape of the unimodal posterior in Figure 1 is considerably simpler than that of the bimodal likelihood, and thus NPE will be easier to converge than NLE. However, this is the opposite for some applications presented in Section 4, where the likelihood is simpler and thus NLE converges much faster than NPE or NRE. In cases where both the posterior and likelihood exhibit complicated shapes, NRE models are a strong option, as they do not require an explicit choice of NDE. Secondly, a good rule of thumb is that neural networks are easier to train when using high-dimensional input to produce low-dimensional output than vice versa. Thus, if $\operatorname{dim}(\bm{x})$ is large (e.g., for image- or graph-like data) then NPE may work better than NLE models. Once again, NRE models are a nice middle-ground as they take in both $\bm{x}$ and $\bm{\theta}$. Note, this heuristic may be changing with the advent of diffusion generative models (Song et al.(2020)Song, Sohl-Dickstein, Kingma, Kumar, Ermon, and Poole), but those are not yet integrated into LtU-ILI. Lastly, it is imperative to consider the downstream applications of the inference. NPE models are much better for cases in which the inference must be repeated for many test $\bm{x}_{\rm o}$ (e.g., estimating morphology for many JWST galaxies), as it can be prohibitively expensive to run MCMC chains for NLE and NRE for many test points. However, if the learned posteriors will simply be injected into a broader hierarchical likelihood sampling (e.g., using cluster masses to constrain cosmology), NLE or NRE would be a more natural choice.
@@
For a quantitative comparison of NPE, NLE, and NRE on machine learning benchmarks, see Lueckmann et al.(2021)Lueckmann, Boelts, Greenberg, Goncalves, and Macke. In any case, the best way to choose a method for a new problem is to try each and perform quantitative comparisons, tests which are made considerably easier with the configuration interface in LtU-ILI.
@@ 2.3 Sequential Learning
Thus far, we have phrased our problem in such a way that we first have a set of pre-run simulations, and only later do we obtain the posterior distribution given our observations. This can be the case when the simulations are particularly expensive (e.g., $N$-body simulations in cosmology, or Earth systems models for climate science e.g., from which Watson-Parris et al.(2021)Watson-Parris, Williams, Deaconu, and Stier built emulators). In these cases, we must be careful about any difference between the proposal prior for the parameters from which the simulations were run (e.g., a Latin hypercube) and the chosen prior for the parameters used in the inference (e.g., a multivariate Gaussian). It may be the case that this proposal prior is significantly wider than the posterior distribution inferred once the real data have been observed, meaning that many of the simulations used in training for NPE/NLE/NRE have been “wasted” as the posterior has almost no support in the region of parameter space occupied by a (potentially significant) fraction of the simulations.
One can instead adopt a multi-round inference approach to improve simulation efficiency, referred to as Sequential NPE/NLE/NRE or SNLE/SNLE/SNRE. In this method, we begin by running a fraction of the total desired simulations with parameters sampled from the broad prior, conducting the first round of training for the NPE/NLE/NRE model. This weak inference is then applied to $\bm{x_{\rm o}}$ to derive a first estimate for the posterior. The algorithm then identifies regions of high posterior density and conducts additional simulations within these areas to produce a refined posterior estimate. This can be repeated until a convergence criterion is met or the total simulation budget is exhausted. This iterative refinement of the posterior around the region of interest can lead to improved posterior estimates for a given experiment. However, it is important to note that these results are deliberately optimized for a singular observed data point. The sampled simulations may not be appropriate for a different experiment, potentially resulting in a less accurate estimate of the posterior distribution if no further simulations are run.
Note that an alternative method for generating parameters of interest using acquisition functions that minimize the expected uncertainty in the posterior density approximation exists. This is based on the Bayesian Optimization for Likelihood-Free Inference (BOLFI) framework (Gutmann et al.(2016)Gutmann, Cor et al.) has also been suggested and applied to cosmological data (Leclercq(2018)), although this is not yet part of LtU-ILI.

### v2

@@ 2.2 Neural Density Estimation
In modern applications, implicit inference is made tractable by using machine learning models to emulate conditional probability distributions, a procedure called Neural Density Estimation (Papamakarios, 2019). As an example, consider a situation wherein we seek to train a mixture density network (MDN; Bishop, 1994) to directly model the conditional distribution $\mathcal{P}(\bm{v}|\bm{u})$ using our training set of data-parameter pairs $(\bm{x},\bm{\theta})\sim\mathcal{D}_{\rm train}$. For this subsection only, we have $(\bm{u},\bm{v})=(\bm{x},\bm{\theta})$ or $(\bm{u},\bm{v})=(\bm{x},\bm{\theta})$ interchangeably. The goal is to build a neural architecture $q_{\bm{w}}(\bm{v}|\bm{u})$ with weights $\bm{w}$ that outputs a probability distribution over $\bm{v}$ which targets the conditional probability $\mathcal{P}(\bm{v}|\bm{u})$. An MDN outputs probability distributions by allowing the neural network to specify the statistical parameters of a mixture density distribution. In the independent Gaussian case, this means the neural network outputs a mean and a variance for each component of the mixture density distribution, i.e.,
@@
where the expectation $\mathbb{E}_{\mathcal{D}_{\rm train}}$ is taken over all data-parameter pairs in the training set $\sim\mathcal{D}_{\rm train}$. For the case of the the posterior distribution ($\bm{v}=\bm{\theta},\bm{u}=\bm{x}$), if the training data is sufficiently diverse and if the MDN is sufficiently flexible, then the minimization of $\mathcal{L}_{\rm MDN}$ over the network weights $\bm{w}$ will converge $q_{\bm{w}}(\bm{\theta}|\bm{x})$ to the true posterior $\mathcal{P}(\bm{\theta}|\bm{x})$ (Papamakarios and Murray, 2016).
The MDN is just one example of an ever-increasing variety of neural density estimators in the literature. Another very popular class of methods is normalizing flows (Papamakarios et al., 2021), which output a flexible conditional distribution defined via learnable, invertible transformations of a base Gaussian distribution. Normalizing flows are often better equipped for handling cases wherein the target distribution is non-Gaussian, but low-dimensional and uni-modal. For any type of neural density estimator, the training procedure always involves minimizing the negative log-probability of training data, as in Equation 3.
@@ 2.2.1 Neural Posterior Estimation (NPE)
The most straightforward method to do this is to train neural networks to directly emulate the posterior distribution, also known as Neural Posterior Estimation (NPE; Papamakarios and Murray, 2016; Greenberg et al., 2019). As in the MDN example in the previous section, the loss function for NPE is simply the negative log-posterior of our learned neural density estimator $\hat{\mathcal{P}}(\bm{\theta}|\bm{x})$, and can be trained using a loss function like that of Papamakarios and Murray (2016):
@@
where our neural posterior $\hat{\mathcal{P}}(\bm{\theta}_{i}|\bm{x}_{i})$ is decomposed into a neural network output $q_{\bm{w}}(\bm{\theta}|\bm{x})$ and a weighting factor taken as the ratio of the assumed prior $p(\bm{\theta})$ to the proposal prior $\tilde{p}(\bm{\theta})$. The proposal prior is defined as the distribution of $\bm{\theta}$ present in the training dataset $\mathcal{D}_{\rm train}$, while the assumed prior is an experimental design choice representing the assumed knowledge of the global distribution of $\bm{\theta}$. In many cases, the assumed and proposal priors are identical, however this distinction is particularly relevant in the case of sequential learning (see section 2.4). Note, that the normalization factor $p(\bm{\theta})/\tilde{p}(\bm{\theta})$ cancels out when optimizing $\mathcal{L}_{\rm NPE}$ over network weights $\bm{w}$, but is still required for constructing the posterior estimator $\hat{\mathcal{P}}(\bm{\theta}|\bm{x})$.
This method has the advantage that it allows for quick evaluation and sampling of the full posterior at inference time. This accelerated emulation of the posterior is often called amortized inference. However, the disadvantage is that it requires knowledge of the analytic form of the proposal prior $\tilde{p}(\bm{\theta})$ from which training data was sampled, which may not be accessible for astrophysical training data (e.g., cluster properties in cosmological simulations). Recently, Vasist et al. (2023) used an NPE strategy to constrain exoplanetary atmospheric models. NPE has also become a known tool in Gravitational-Wave astronomy where analytical Compact Binary Coalescent (CBC) waveforms play the role of the forward model (e.g., with DINGO from Dax et al. 2021a, and the constraints on GW150914 by Crisostomi et al. 2023, among many others).
@@ 2.2.2 Neural Likelihood Estimation (NLE)
An alternative method which circumvents the issue of fixed priors is to only fit for the likelihood, $\mathcal{P}\left(\bm{x}|\bm{\theta}\right)$, via Neural Likelihood Estimation (NLE; Alsing et al., 2018, 2019; Papamakarios et al., 2019). The loss function for NLE notably does not include the assumed or proposal prior, and simply maximizes the global log-likelihood,
@@
wherein we emphasize the swapped arguments of $q$ as compared to Equation 4. Multiplying a learned likelihood with an assumed prior results in a proxy which is proportional to the posterior, i.e., $\hat{\mathcal{P}}(\bm{\theta}|\bm{x})\propto q_{\bm{w}}(\bm{x}|\bm{\theta})p(\bm{\theta})$. Generating samples from the posterior is then simply a matter of using this proxy in running Markov chain Monte Carlo (MCMC; Robert et al., 1999) or Variational Inference (Blei et al., 2017). For example, in Metropolis-Hastings MCMC (Robert et al., 2004), the learned likelihood and assumed prior can be used to calculate the acceptance probability of a transition $\bm{\theta}_{t}\rightarrow\bm{\theta}^{\prime}$ without knowledge of the full normalized posterior, i.e.,
@@
where $q(\bm{\theta}^{\prime}|\bm{\theta}_{t})$ is the proposal mechanism. Note, the process of obtaining a posterior is identical to the explicit likelihood-based approach, but instead of an analytic likelihood function, NLE has essentially a “black box” function instead. The advantage of this method is that one only needs to perform training once for a given model, at the cost of additional sampling once data becomes available. For instance, Jeffrey et al. (2021) fit a density model for the likelihood of compressed summaries from DES SV weak lensing maps. The Madminer tool (Brehmer et al., 2020) for particle physics notably allows for NLE using ILI tools.
@@ 2.2.3 Neural Ratio Estimation (NRE)
As in NLE, Neural Ratio Estimation (NRE; Hermans et al., 2020) targets a proxy that is proportional to the posterior. Instead of outputting a conditional density estimate for the likelihood, NRE models instead target the likelihood ratio,
@@
equal to the ratio of the likelihood at a given point in parameter space to the likelihood at a reference point, $\bm{\theta}_{0}$. This formulation is convenient because it can be folded into an acceptance ratio like Equation 6 and also be cast as a classification problem (Cranmer et al., 2015). The output of a classification network, $d_{\bm{w}}(\bm{x},\bm{\theta})\in[0,1]$, can be considered equivalent to a likelihood ratio as
@@
wherein we associate a positive classification with assigning a data $\bm{x}$ to the correct $\bm{\theta}$, and a negative classification with those assigning data to a $\bm{\theta}^{\prime}$ sampled from the prior. The main benefit of NRE over NPE and NLE is that one need not specify an approximating distribution such as MDNs or normalizing flows to emulate a posterior. The complexity is purely limited in terms of the depth and design of the neural architecture. Among others, Cole et al. (2022) and Karchev et al. (2023) applied a variation of NRE (namely Truncated Marginal NRE, Miller et al. 2021) to CMB data and Supernova data, respectively. Additionally, Delaunoy et al. (2020) and Bhardwaj et al. (2023) use NRE for Gravitational Wave parameter inference.
2.3 Choosing an Estimator
@@
The optimal choice of NPE, NLE, or NRE is highly dependent on the properties on the underlying problem, the dimensionalities of $\bm{x}$ and $\bm{\theta}$, and the intended application. First, the ease of learning the shape of a likelihood or posterior distribution can vary dramatically between different physical problems. For example, the shape of the unimodal posterior in Figure 1 is considerably simpler than that of the bimodal likelihood, and thus NPE will be easier to converge than NLE. However, this is the opposite for some applications presented in Section 4, where the likelihood is simpler and thus NLE converges much faster than NPE or NRE. In cases where both the posterior and likelihood exhibit complicated shapes, NRE models are a strong option, as they do not require an explicit choice of NDE.
Another strong rule of thumb is that neural networks are easier to train when using high-dimensional input to produce low-dimensional output than vice versa. Thus, if $\operatorname{dim}(\bm{x})$ is large (e.g., for image- or graph-like data) then NPE may work better than NLE models. Alternatively, for high-dimensional posterior inference, i.e., when $\operatorname{dim}(\bm{\theta})$ is large, NLE often performs better than NPE, especially when strong degeneracies exist between parameters. The NRE implementations in LtU-ILI also struggle in this regime, as evidence suggests that NRE methods require truncation or marginalization to resolve regions of high posterior density (Miller et al., 2021, 2022). This heuristic may change with the advent of extremely flexible diffusion architectures (e.g., Song et al., 2020), but such models are not yet integrated into LtU-ILI.
Lastly, it is imperative to consider the downstream applications of the inference. NPE models are much better for cases in which the inference must be repeated for many tests $\bm{x}_{\rm o}$ (e.g., estimating morphology for many JWST galaxies), as it can be prohibitively expensive to run MCMC chains for NLE and NRE for many test points. However, if the learned posteriors are going to be injected into a broader hierarchical likelihood sampling (e.g., using cluster masses to constrain cosmology), NLE or NRE will be a more natural choice.
@@
For a quantitative comparison of NPE, NLE, and NRE on machine learning benchmarks, see Lueckmann et al. (2021). In any case, the best way to choose a method for a new problem is to try each and perform quantitative comparisons, tests which are made considerably easier with the configuration interface in LtU-ILI.
@@ 2.4 Sequential Learning
Thus far, we have phrased our problem in such a way that we first have a set of pre-run simulations, and only later do we obtain the posterior distribution given our observations. This can be the case when the simulations are particularly expensive (e.g., $N$-body simulations in cosmology, or Earth systems models for climate science e.g., from which Watson-Parris et al. 2021 built emulators). In these cases, we must be careful about any difference between the proposal prior for the parameters from which the simulations were run (e.g., a Latin hypercube) and the chosen prior for the parameters used in the inference (e.g., a multivariate Gaussian). It may be the case that this proposal prior is significantly wider than the posterior distribution inferred once the real data have been observed, meaning that many of the simulations used in training for NPE/NLE/NRE have been “wasted” as the posterior has almost no support in the region of parameter space occupied by a (potentially significant) fraction of the simulations.
One can instead adopt a multi-round inference approach to improve simulation efficiency, referred to as Sequential NPE/NLE/NRE or SNPE/SNLE/SNRE. In this method, we begin by running a fraction of the total desired simulations with parameters sampled from the broad prior $p(\bm{\theta})=\tilde{p}(\bm{\theta})$, conducting the first round of training for the NPE/NLE/NRE model. This weak inference is then applied to $\bm{x_{\rm o}}$ to derive a first estimate for the posterior. For instance, in the case of NPE, this is equivalent to minimizing Equation 4 with $p(\bm{\theta})=\tilde{p}(\bm{\theta})$ during the first round, which gives a model $q_{\bm{w}}(\bm{\theta}|\bm{x})$, and then retraining the model with new simulations drawn from $\tilde{p}(\bm{\theta})=q_{\bm{w}}(\bm{\theta}|\bm{x}=\bm{x}_{\rm o})$ for the second round, and so on. The algorithm (e.g., Greenberg et al., 2019) then identifies regions of high posterior density around the observation $\bm{x}_{\rm o}$ and conducts additional simulations within these areas to produce a refined posterior estimate. This can be repeated until a convergence criterion is met or the total simulation budget is exhausted. This iterative refinement of the posterior around the region of interest can lead to improved posterior estimates for a given experiment. However, it is important to note that these results are deliberately optimized for a singular observed data point. The sampled simulations may not be appropriate for a different experiment, potentially resulting in a less accurate estimate of the posterior distribution if no further simulations are run.
Note that an alternative method for generating parameters of interest using acquisition functions that minimize the expected uncertainty in the posterior density approximation exists. This is based on the Bayesian Optimization for Likelihood-Free Inference (BOLFI) framework (Gutmann et al., 2016) has also been suggested and applied to cosmological data (Leclercq, 2018), although this is not yet part of LtU-ILI.